Assume Poland’s currency (the zloty) is worth $.20 and the Japanese yen is worth $.01. What is the cross rate of the zloty with respect to yen? That is, how many yen equal a zloty?
Answer #1

Cross Rate of Zloty to Yen = 1/Dollar to Zloty * Dollar to Yen =1/0.20*0.01 = 0.05

1 zloty equals =1/0.05 = 20
Answer #2

To find the cross rate of the zloty with respect to yen, we need to convert both currencies to a common currency, such as the US dollar, and then compare their values.

Given that 1 zloty is worth $.20, we can say:

1 zloty = $0.20

Similarly, given that 1 yen is worth $.01, we can say:

1 yen = $0.01

To find the cross rate, we need to express the value of 1 zloty in terms of yen. We can do this by first converting the value of 1 zloty from dollars to yen:

1 zloty = $0.20 = 0.20 / 0.01 yen = 20 yen

Therefore, the cross rate of the zloty with respect to yen is 20 yen per zloty.


answered by: Hydra Master